OpenStudy (anonymous):

systems of linear equations and word problems half of isabel's age added to 1/3 of frank's age is 11 years. 6 years from now, the sum of their ages will be 40. how old is each now?

OpenStudy (anonymous):

\[(1/2)i+(1/3)f= 11 \]\[i+f+6=40\]\[i=40-6-f\]\[i=34-f\]\ now substitute "i" into the first equation to find the value of "f"\[(1/2)(34-f)+(1/3)f=11\]\ then solve for f (you can do this!)... once you know "f" then substitute that value in and solve for "i"

OpenStudy (anonymous):

thank you so much :)